Friday the 13th day shift summary (08/13/99)

Ashkan_Alavi-Harati (
Fri, 13 Aug 1999 15:36:33 -0500

Shift Summary for day shift, 13/Aug/99

On Shift: U. Erkarslan, Ashkan

Hours of beam available at all intensities: 6.5/8.5
Hours of beam available at nominal intensity: 6.0/8.5
Hours of E832 data (special runs not included):6.5/8.5
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:
Various accelerator problems.

Runs ended this shift:

Run 13222 - in progress - BEAM_832

Tapes brought to FCC this shift:
(including tapes for online split)
KQC255-257 KQB090

Accesses during shift: none

Overview of shift:
Smooth E832 datataking for most of the shift.

Description of Shift:

run spill time comment
14222 127 07:00 Take over the ongoing run from the owl shift.
193 08:30 No beam for spills 193-198.
264 10:05 Beam is gone to fix a MI power supply.
282 10:28 Beam is back.
286 10:35 No beam again. MI vacuum problem.
305 11:00 Beam is back.
309 11:07 MI-8GeV line PS problem.
324 11:26 Beam is back but the intensity will vary while tuning.
339 11:46 beam reaches the nominal intensity.
405 13:15 Noticed beam is completely off the target in vertical
direction, called MCR.
407 13:19 beam back on but MCR says one of the power supplies
in our SY is not ramping and they may have to replace a
449 14:10 tapes are full, run ends.

14223 1 14:20 Start the run, beam was interrupted for 6 spills while
we were changing tapes. So we didn't lose any time!
14:30 fix_disp screens 5 and 6.
35 15:00 Tevatron injection tuning, beam intermitent until 14:00

Follow up Notes on Previous Shift

Items to Follow up on Next Shift

Please do a dart fresh start next time you start a run, since we had to
fix two of the displays.